An electrical firm manufactures light bulbs that have a length of life that is approximately normally distributed, with mean equal to 810 hours and a standard deviation of 74 hours. Find the probability that a random samples of 36 bulbs will have an average life of greater than 803 hours


1
Expert's answer
2022-02-10T17:53:30-0500

Let "X=" an average life: "X\sim N(\mu, \sigma^2/n)."

Given "\mu=810\ h, \sigma=74\ h, n=36."

"P(X>803)=1-P(X\le 803)"

"=1-P(Z\le \dfrac{803-810}{74/\sqrt{36}})"

"\approx 1-P(Z\le-0.567568)"

"\approx0.714836"